OpenStudy (anonymous):

You have a 5 -question multiple- choice test. Each question has four choices. you don't know any of then answers. what is the experimental probability that you will guess exactly three out of five questions correctly ?

OpenStudy (anonymous):

The experimental? You'd need to try it yourself. The theoretical? We have a \(\frac{1}{4}\) chance in getting it right, and three questions to get right, thus: \[ \frac{1}{4}\cdot\frac{1}{4}\cdot\frac{1}{4}=\frac{1}{64} \]Which is our theoretical probability of getting all three questions correct.